OpenStudy (anonymous):

Factor completely: 2x2 - 4x - 48

OpenStudy (anonymous):

2(x^2 - 2x - 24) 2(x - 6)(x + 4)

OpenStudy (anonymous):

2x^2-4x-48=0 x^2-2x-48=0 x^2-6x+4x-24=0 x(x-6)+4(x-6)=0 (x-6)(x+4)=0 which x=-4,6